# Characterisation of Open Quantum System Dynamics based on Information Back-flow

### v1

@@ IV Information flow and distinguishability
The idea of information flow was introduced as an interpretation of distinguishability in the BLP definition. If the distinguishability is invariant between all pairs, then this shows no information exchange, indicating that the dynamics is unitary. On the other hand, if the distance monotonically decreases, the distinguishability decreases, suggesting that the information should be transferred away from the system.
This intuition was justified by defining the internal information ($I_{int}$) between the system states by BLP and GBLP distances, which can be accessed by the measurement of the system alone Breuer et al. (2016). The external information ($I_{ext}$) which cannot be accessed alone from the system lies in correlation or/and within environment states and is defined by subtracting from total information ($I_{total}$), which is from the combined state of the system-environment. So
$I_{int}^{t}(\rho_{S}^{1},\rho_{S}^{2})+I_{ext}^{t}=I_{total}^{t}(\rho^{1},\rho^{2}),$ (10)
where $I_{ext}=I_{total}-I_{int}$ and $\rho_{S}$, $\rho$ are system and combined system-environment state respectively. Since total information is constant owing to the unitarity of the dynamics, $I_{total}^{t}=I_{total}^{0}$. All arguments of information forward-flow or back-flow are justified by Eq. (10). We now consider how well the shift behaviour in a non-unital noise connects with the distinguishability condition.
IV.1 False positive test of distinguishability
To give an example of a case where the GBLP condition can be misleading, consider the dynamical map with $T=\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&0\\ \end{array}\right)$, which takes all the states to the centre of the Bloch sphere, followed by shift $\vec{c}=\left(\begin{array}[]{c}0\\ 0\\ c\end{array}\right)$ which takes that centre to $c$. If we apply the GBLP measure, the distance will be $\lvert(p-q)\rvert$ or $\lvert(p-q)c\rvert$ depending on the value of $c$, which is non-zero. Though the states are identical, this distance gives the appearance of distinguishing them because it has a label of biased probabilities, which reflects the preparation information rather than information back-flow.
The divisible dynamics mentioned in theorem 2 where GBLP is necessary will always give false positives. That can specifically be seen in the example shown in II.2.
IV.2 False negative test of distinguishability
The GBLP definition does not follow distinguishability and information flow for all pairs. There is a possibility states can evolve and come closer or go far, but the information flow is zero, which misguides the interpretation of distinguishability and information flow.
@@ IV.2 False negative test of distinguishability
$T=\left(\begin{array}[]{ccc}\exp^{-\gamma t}&0&0\\ 0&\exp^{-\gamma t}&0\\ 0&0&\exp^{-\gamma t}\end{array}\right)$ (11)
@@
According to Eq. (10), the BLP definition suggests the information outflow for all pairs for all time. In light of Eq. (9), the generalized distance is constant for Bloch radius $r\leq 0.5$ in the considered example for $\gamma=0.1$ and $p=0.25$. Thus, GBLP will show no outflow or distinguishability change for certain pairs for corresponding probabilities, as shown in Fig (2). This goes against our expectation that the absence of information flow should only be true for unitary dynamics or the pairs where their sub-spaces do not evolve. This suggests that the generalized definition of distance does not capture the information flow for all pairs as required by Eq. 10.
@@
Figure 2: Generalised distance is plotted as a function of time for different Bloch radii $r$ with two diametrically opposite points for an isometrically decaying dynamical map (Eq. 11). The decay parameter $\gamma$ and the probability of preparing states $p$ are 0.1 and 0.25, respectively. Distance and time are in arbitrary units.
@@
Figure 3: Generalised distance is plotted as a function of time for different pairs of diametrically opposite points on the surface of Bloch sphere for a spin interacting with another spin (Eq. 12). The dynamical parameter $\omega$ and the probability of preparing states $p$ are 1.25 and 0.25, respectively. Distance and time are in arbitrary units.
@@
This issue is not restricted to BLP Markovian dynamics. For BLP non-Markovian dynamics, the domain exists, showing no outflow or inflow, even though states manifestly evolved during that period. To show this, two interacting spins are considered; one spin is considered a system and the other an environment. The dynamical map is given by:
$T=\left(\begin{array}[]{ccc}\cos{\omega t}&0&0\\ 0&\cos{\omega t}&0\\ 0&0&1\\ \end{array}\right)$ (12)
@@
This non-invertible process is trivially non-Markovian and has an evident information back-flow. Yet, for certain pairs of states, it can evolve their distance from a regime of $\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert$ to one of $\lvert\vec{w}(t)\rvert<\lvert p-q\rvert$, which in light of Eq. (8), can cause the GBLP measure to flatten out. The non-changing phase is shown in Fig (3) for $\omega=1.25$ and $p=0.25$, which shows no information flow from or towards the system. We can notice in this example that pairs taken on the z-axis will not evolve, which is consistent for both BLP and GBLP measures.
These examples suggest that GBLP is a measure of contracting maps related to P-divisibility. This does not capture the idea of information flow for all cases.
@@ V Discussion and Conclusion
The generalised BLP measure was needed to characterize non-Markovianity because BLP did not capture aspects of the non-unital noise. In specific, it was shown that GBLP has a direct connection with P-divisibility. We have demonstrated that GBLP and BLP are equivalent witnesses of non-Markovianity for unital dynamics. For non-unital cases, where the Bloch sphere is contracting, BLP does not help decide the non-Markovianity. In those cases, GBLP shows that distance can increase with time solely because of translation, which is associated with information back-flow. Here, we have indicated certain situations that undermine this association. The GBLP condition can yield a false positive result as an artifact of initial biased preparation. Conversely, we also indicate instances where it yields a false negative result when some initial pairs do not change the generalized distance despite the evident effect of the environment on their subsystem dynamics. Thus, the GBLP condition is only correct for capturing the divisibility properties, but not information flow for all cases.
@@ Acknowledgments
VP acknowledges the financial assistance of the Indian Science and Engineering Research Board (SERB) through grant CRG/2022/008345.

### v2

@@ IV Information flow and distinguishability
The idea of information flow was introduced as an interpretation of distinguishability in the BLP definition. If the distinguishability is invariant between all pairs, then this indicates no information exchange, suggesting that the dynamics is unitary. On the other hand, if the distance monotonically decreases, the distinguishability decreases, suggesting that the information should be transferred away from the system.
This intuition was justified by defining the internal information ($I_{int}$) between system states using BLP and GBLP distances, which can be accessed through the measurement of the system alone Breuer et al. (2016). The external information ($I_{ext}$), which cannot be accessed alone from the system, lies in correlation and/or within environment states and is defined by subtracting from total information ($I_{\rm total}$), which is from the combined state of the system-environment. So
$I_{\rm int}^{t}(\rho_{S}^{1},\rho_{S}^{2})+I_{\rm ext}^{t}=I_{\rm total}^{t}(\rho^{1},\rho^{2}),$ (11)
where $I_{\rm ext}=I_{\rm total}-I_{\rm int}$ and $\rho_{S}$, $\rho$ are system and combined system-environment state respectively. Since total information is constant owing to the unitarity of the dynamics, $I_{\rm total}^{t}=I_{\rm total}^{0}$. All information forward-flow or back-flow arguments are justified by Eq. (11). This suggests that information flow should be associated with individual pairs of states, unlike properties such as P-divisibility or CP-divisibility, which characterise a map as a whole. We now consider how tight or faithful a witness the GTD criterion is of information flow in this sense. Below, we will provide specific instances to argue that it is neither a tight nor a faithful witness.
IV.1 False negative test of distinguishability
We now demonstrate that the GTD criterion is not a tight witness of distinguishability when applied to specific pairs of states. The GBLP definition does not follow distinguishability and information flow for all pairs at all times. There is a possibility that states can evolve and come closer or go far, but the information flow is indicated to be zero.
@@ IV.1 False negative test of distinguishability
$T(t)=\left(\begin{array}[]{ccc}\exp^{-\gamma t}&0&0\\ 0&\exp^{-\gamma t}&0\\ 0&0&\exp^{-\gamma t}\end{array}\right)$ (12)
@@
According to Eq. (11), the BLP definition suggests the information outflow for all pairs for all time. In light of Eq. (10), the generalized distance is constant for Bloch radius $r\leq 0.5$ in the considered example for $\gamma=0.1$ and $p=0.25$. Thus, GBLP will show no outflow or distinguishability change for certain pairs for corresponding probabilities, as shown in Fig (2). This contradicts our expectation that the absence of information flow should only be true for unitary dynamics or pairs where their subspaces do not evolve. This suggests that the generalized definition of distance does not capture the information flow for all pairs as required by Eq. 11.
@@
Figure 2: generalized distance is plotted as a function of time for different Bloch radii $r$ with two diametrically opposite points for an isometrically decaying dynamical map (Eq. 12). The decay parameter $\gamma$ and the probability of preparing states $p$ are 0.1 and 0.25, respectively. Distance and time are in arbitrary units.
@@
Figure 3: Generalized distance is plotted as a function of time for different pairs of diametrically opposite points on the surface of the Bloch sphere for a spin interacting with another spin (Eq. 13). The dynamical parameter $\omega$ and the probability of preparing states $p$ are 1.25 and 0.25, respectively. Distance and time are in arbitrary units.
@@
This issue is not restricted to BLP Markovian dynamics. For BLP non-Markovian dynamics, the domain exists, showing no outflow or inflow, even though states manifestly evolved during that period. To illustrate this, two interacting spins are considered, with one spin serving as the system and the other as the environment. The dynamical map is given by:
$T(t)=\left(\begin{array}[]{ccc}\cos{\omega t}&0&0\\ 0&\cos{\omega t}&0\\ 0&0&1\\ \end{array}\right)$ (13)
@@
This non-invertible process is trivially non-Markovian and exhibits evident information backflow. Yet, for certain pairs of states, it can evolve their distance from a regime of $\lvert\vec{w}(t)\rvert\geq\lvert p-q\rvert$ to one of $\lvert\vec{w}(t)\rvert<\lvert p-q\rvert$, which, in light of Eq. (9), can cause the GBLP measure to flatten out. The non-changing phase is shown in Fig (3) for $\omega=1.25$ and $p=0.25$, which shows no information flow from or towards the system. We can observe in this example that pairs along the z-axis do not evolve, which is consistent for both BLP and GBLP measures.
These examples suggest that GBLP is a measure of contracting maps related to P-divisibility. This does not capture the idea of information flow for all cases at all times. Even in our examples, we see that there are other time intervals where GBLP faithfully captures information outflow or backflow (the non-flat regimes of in Figure 3), and other states where GBLP faithfully captures information outflow or backflow at all times. We stress that our result doesn’t undermine the optimality of GBLP as a witness of the non-P-divisibility of a map but rather highlights that it isn’t a faithful witness of information backflow in the context of individual pairs of states.
IV.2 False positive test of distinguishability
More worryingly, we now demonstrate that the GTD criterion can be unfaithful in the sense we discussed when applied to specific pairs of states. To give an example of a case where the GBLP condition can be misleading, consider the composition of two channels as $\mathcal{E}_{t,0}=\mathcal{B}_{t,\tau}\circ\mathcal{A}_{\tau,0}$ where $\mathcal{A}$ acting till time $\tau$ starting from zero is a CPTP depolarising channel $\mathcal{A}(\rho)=(1-p)\rho+p\frac{\mathbb{I}}{2}$ corresponds to the unital channel with matrix form given by $\mathcal{A}(\tau)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1-p(\tau)&0&0\\ 0&0&1-p(\tau)&0\\ 0&0&0&1-p(\tau)\\ \end{array}\right)$ and $\mathcal{B}$ is the shift channel given by $\mathcal{B}(t,\tau)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ c&0&0&0\\ \end{array}\right)$. Here, the depolarising probability is $p(0)=0$ and $p(\tau)=1$. The shift is assumed to be sufficiently small not to leave the Bloch ball. More generally, suppose the shift dynamics is given by an increase in Bloch vector length given by $c(1-|\vec{r}|)$, which ensures that positivity is preserved. At time $\tau$, all initial states will go to a completely mixed state. The GBLP measure evaluates the distance to zero at time $\tau$, and after the shift channel, the distance will be $\lvert(p-q)\rvert$ at time $t$ irrespective of the value of $c$. Though the states are identical at time $\tau$, this non-zero distance after the shift gives the appearance of distinguishing them because it has a label of biased probabilities, which reflects the preparation information rather than information back-flow from the environment.
The divisible dynamics mentioned in Theorem 2, where GBLP is necessary, will necessarily give false positives. This can be specifically seen in the example shown in II.2.
@@ V Discussion and Conclusion
The generalized BLP measure was needed to characterise the non-Markovianity of maps because BLP did not capture aspects of the non-unital noise. Specifically, it was shown that GBLP has a direct connection with the P-divisibility of maps. We have demonstrated that GBLP and BLP are equivalent witnesses of non-Markovianity for unital dynamics. For non-unital cases where the Bloch sphere is contracting, BLP does not help decide the non-Markovianity of maps. In those cases, GBLP shows that distance can increase solely with time due to the translation term associated with information backflow. Here, we have argued that for GBLP to be a meaningful (faithful) witness of information backflow (distinguishability), GBLP should apply to any individual pair of states rather than to just the map as a whole. We have indicated certain situations that undermine the status of GBLP as a faithful witness of information backflow (distinguishability). Specifically, the GBLP condition can produce a false negative result when some initial pairs do not change the generalized distance despite the evident effect of the environment on their subsystem dynamics. This entails that the condition is not a tight witness of information flow, weakening its universal applicability over the state space. More seriously, we also indicate instances where it yields a false positive result as an artifact of the initial non-equiprobable preparation of two states and not because of the system-environment interaction. Thus, while the GBLP condition optimally witnesses the divisibility properties of the map as a whole, it fails to witness information flow for all cases faithfully. In general, this result highlights that the distinguishability increases, which suggests the information backflow for one pair is enough to decide the non-Markovianity of the map as a whole, but that does not guarantee that the condition serves as a faithful witness in the context of distinguishability and information flow analysis, because that is meaningful only when applicable to all pairs. This result highlights that any definition used to witness non-Markovianity through at least one instance does not automatically become a witness of distinguishability or information backflow. Here, the GBLP case for a qubit, as an example, is considered to highlight this difference; other specific definitions and higher dimensions can be explored in the future. What kind of information non-unital dynamics possess within their translation part should also be explored to understand the information aspect of non-Markovianity for any dynamics.
@@ Acknowledgments
VP acknowledges the financial assistance of the Anusandhan National Research Foundation (ANRF) through grant CRG/2022/008345.